# arXiv:2603.00942v1[math.DG] 01 Mar 2026

### v1

@@ 2 Preliminaries and some useful facts
and so
$\sqrt{|g|}=\det\mathbb{A}(s,\xi).$
Therefore, by applying (2.2) and (2.3), it follows that the volume $\mathrm{vol}(B(q,r))$ of the geodesic ball $B(q,r)$ on $M^{n}$ is given by
@@
One has $\ell(q)=\max_{\xi\in S_{q}M^{n}}d_{\xi}$ (see page 704 of [31, Section 2] for the explanation). Clearly,
@@
As pointed out in [31, Remark 2.4] or [47, Remark 2.5], since the radial distance is $r(x)=d_{M^{n}}(x,q)$ for $x=\gamma_{\xi}(s)=\exp_{q}(s\xi)$, the parameter $s$ may be seen as the argument of the continuous function $\kappa:[0,l)\rightarrow\mathbb{R}$ in Definition 2.2 or Definition 2.3. Therefore, one has $\frac{d}{ds}|_{x}=\nabla r(x)=v_{x}$, which implies (2.12) and (2.13) become
@@
Similarly, one can define the concept that the radial Ricci/sectional curvature has an upper/lower bound (with respect to some point) which is given by a continuous function of the radial distance parameter. By the way, the concept that the radial (Ricci or sectional) curvature is bounded at some point has been improved to a more general setting, i.e. the integral radial (Ricci or sectional) curvature is bounded with respect to some point – see [51, Definitions 2.1 and 2.2]. In fact, the quantities defined in [51, Definitions 2.1 and 2.2] can be used to measure the deviation of the radial (Ricci or sectional) curvature (with respect to some point $q\in M^{n}$) from a prescribed continuous function $(n-1)\kappa(r(x))$ or $\kappa(r(x))$, where $r(x)=d_{M^{n}}(x,q)$ denotes the intrinsic distance to $q$ on the Riemannian $n$-manifold $M^{n}$.
Now, we wish to point out an important truth, that is, for a prescribed $n$-dimensional complete manifold $M^{n}$, one can always construct the optimal continuous functions $\kappa_{\pm}(q,s)$ with respect to a point $q\in M^{n}$, satisfying Definitions 2.2 and 2.3 respectively. This truth has already been shown in our previous works [31, p. 706], [47, pp. 378-379]. However, for convenience to readers, we wish to repeat the construction here. Recall that for $\xi\in S_{q}^{n-1}\subseteq T_{q}M^{n}$, $\gamma_{\xi}(s)=\exp_{q}(s\xi)$ and its derivative $\gamma^{\prime}_{\xi}(s)$ depend smoothly on the variables $(s,\xi)$. Set $\mathbb{D}_{q}:=\{(s,\xi)\in[0,\infty)\times S_{q}^{n-1}|0\leq s<d_{\xi}\}$ with closure $\overline{\mathbb{D}}_{q}:=\{(s,\xi)\in[0,\infty)\times S_{q}^{n-1}|0\leq s\leq d_{\xi}\}$. Then we can define
@@
As explained in [31, pp. 706-707] or [47, p. 379], by using the properties of warped product manifolds, one knows that the radial sectional curvature, and the radial component of the Ricci tensor of the spherically symmetric manifold $0,l)\times_{f(s)}\mathbb{S}^{n-1}$ with the base point $q$ are given by
@@
Clearly, if $r=0$ and $R=r_{0}$, then the annulus $A^{\Gamma}_{0,r_{0}}(q)$ would be exactly the geodesic ball $B(q,r_{0})$ on the given Riemannian $n$-manifold $M^{n}$. In [49, Theorem 3.2] (resp., [49, Theorem 4.3]), we have successfully obtained a relative volume comparison for the annulus $A^{\Gamma}_{r,R}(q)$ on a complete Riemannian $n$-manifold $M^{n}$ having a radial Ricci curvature lower bound $(n-1)f^{\prime\prime}(s)/f(s)$ with respect to the point $q\in M^{n}$ (resp., a radial sectional curvature upper bound $f^{\prime\prime}(s)/f(s)$ with respect to the point $q$). Moreover, those two volume comparison results would immediately imply our Bishop-type volume comparisons (i.e. Theorems 2.6 and 2.7 here) by suitably choosing radius parameters for annuluses – see [49, Corollaries 3.4 and 4.4] for details. In 2023, we attempted to improve our previous volume comparison results mentioned here to the setting that the given complete Riemannian $n$-manifold has integral radial curvature bounds (i.e. a more weaker curvature assumption), and this attempt was successful – see [51, Section 3] for details.
@@
In [29, Section 3], we have given several lower bounds for the first eigenvalue of weighted $p$-Laplacian on submanifolds with locally bounded weighted mean curvature.
@@ 3 Weighted heat kernel comparison theorems
This question is crucial for the eigenvalue problem (1.11) and the answer is affirmative. In fact, as mentioned in A. Grigor’yan’s survey [36, Section 2], by suitably making adjustments to those related arguments in e.g. [18, 26, 32, 60], it is not hard to get the following result.
@@
([36, Section 2]) The operator $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a self-adjoint nonpositive definite operator in $L^{2}_{\phi}$, where the space $W^{2,2}_{0,\phi}(\Omega)$ is defined as $W^{2,2}_{0,\phi}(\Omega):=\{\rho\in W^{1,2}_{0,\phi}(\Omega)|\Delta_{\phi}\rho\in L^{2}_{\phi}(\Omega)\}$. Moreover, the operator $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a unique self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$ with the domain in $W^{1,2}_{0,\phi}(\Omega)$. If in addition $\Omega$ is geodesically complete, then the operator $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$ is essentially self-adjoint, i.e., $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a unique self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$.
Once we have Theorem 3.1 for the extension of the operator $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$. The spectrum of the eigenvalue problem (1.11) is actually the spectral structure of the self-adjoint operator $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ in $L^{2}_{\phi}(\Omega)$, which is only discrete and can be deduced from the compactness of the resolvent $\left(-\Delta_{\phi}+I\right)^{-1}$, together with the spectral theory of compact self-adjoint differential operators. See e.g. [36, Subsection 2.2] for details. Combining with the variational principle, one can get two basic facts for the eigenvalue problem (1.11) listed in Section 1.
As explained in [36, Section 3], once the self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$ has been set up as in Theorem 3.1, the heat semigroup $e^{t\Delta_{\phi}}$ related to the Witten-Laplacian $\Delta_{\phi}$ can be well-defined. Then one can get the following properties for the weighted heat kernel by using the tool of heat semigroup.
@@
(see e.g. [36, Theorem 3.3 in Section 3]) Let $M$ be a complete Riemannian manifold and $\phi$ be a real-valued smooth function on $M$. Denote by $\mathbb{R}^{+}$ the set of all positive real numbers. Then there exists a weighted heat kernel $H^{\phi}(x,y,t)\in C^{\infty}(M\times M\times\mathbb{R}^{+})$ such that
@@
It is easy to see that (1)-(2) in Lemma 3.2 have close relation with the existence of the weight heat kernel. In [36, Theorem 3.3 in Section 3], the property (3) is the symmetry of the weight heat kernel, while the property (4) is the semigroup identity of the weight heat kernel. In [36, Section 3], A. Grigor’yan mentioned at least two approaches to prove the existence of the weight heat kernel — the classical one is the usage of a parametrix of the heat equation, and the other one is the usage of eigenfunction expansion formula in an exhaustion of relatively compact open sets $\{\Omega_{i}\}_{i=1}^{\infty}$ of $M$. For details, one can check e.g. [10, 15, 18, 26, 60] as references.
@@
It is easy to check that $\mathcal{L}^{\phi}$ is uniformly parabolic, and then the assertion in Theorem 3.1 directly follows from the standard theory of second-order parabolic partial differential equations (see e.g. [15, p. 180] or [30, Chapter 7]). ∎
@@
If the potential function $\phi$ has Property 1, then the weighted heat kernel $H_{+}^{\phi}$ (resp, $H^{\phi}_{-}$) on a geodesic ball $\mathscr{B}_{n}(q^{+},\cdot)$ (resp, $\mathscr{B}_{n}(q^{-},\cdot)$) of the model space $M^{+}$ (resp, $M^{-}$) depends on the space variable $r_{1}$ (resp., $r_{2}$) can be observed by the decomposition of the Witten-Laplacian on $\mathscr{B}_{n}(q^{+},\cdot)$ (resp, $\mathscr{B}_{n}(q^{-},\cdot)$) in geodesic spherical coordinates at $q^{+}$ (resp, $q^{-}$) — see the $3^{\rm{rd}}$ line in the proof of Lemma 6.1 for the detailed expression. The monotonicity of $H_{+}^{\phi}$ (resp, $H^{\phi}_{-}$) with respect to the time variable can be obtained by using an almost the same argument to that in the proof of [18, Lemma 2.3]. But there are two differences we wish to point out as follows:
@@
    where $a_{i}(r_{1})$ are smooth functions on $M^{+}$. This directly implies the expression (2.11) in [18]. In the weighted case, [36, Theorem 3.9] tells us that for sufficiently small $0<r_{1},t<\epsilon$, there exists a smooth positive function $u(r_{1})$ such that the weighted heat kernel $H_{+}^{\phi}(r_{1},t)$ satisfies
@@
    which leads to the truth that the analysis process in the proof of [18, Lemma 2.3] before the equation (2.20) can be transplanted to the non-weighted case without any big obstacle. Same story happens to $H^{\phi}_{-}(r_{2},t)$. By the way, a detailed proof of the above asymptotic expansion for the weighted heat kernel can be found in [9, 34].
@@
We wish to use a similar argument to that in the proof of our previous work [47, Theorem 6.6].
@@
$\qquad H^{\phi}(q,y,t)-H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)$
@@
$\geq 0,$ (3.5)
where the fact $d\theta(s,\xi)/ds=(n-1)\left[J(s,\xi)/f(s)\right]^{n-2}\left[J(s,\xi)/f(s)\right]^{\prime}\leq 0$ shown in Theorem 2.6 has also been used. This directly implies $H^{\phi}(q,y,t)\geq H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)$, which is exactly the weighted heat kernel comparison (1.7). When the equality in (1.7) holds at some $(y_{0},t_{0})\in B(q,r_{0})\times(0,\infty)$, by Theorem 3.4 one has $H^{\phi}(q,y,t)=H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)=H^{\phi}(q,y_{0},t_{0})$ on $B(q,r_{0})\times[0,t_{0}]$. Together with (3), we know that
@@ 4 Applications in spectral geometry
Theorem 3.1 tells us that if $\Omega\subseteq M^{n}$ is geodesically complete, with $M^{n}$ a given complete Riemannian $n$-manifold, then $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a unique self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$, and this fact can be used to prove that the eigenvalue problem (1.11) only has discrete spectrum, which consists of eigenvalues given in the sequence (1.12). Let $\psi_{i}$ be an eigenfunction corresponding to the $i$-th Dirichlet eigenvalue $\lambda_{i,\phi}(\Omega)$ such that $\{\psi_{i}\}_{i=1}^{\infty}$ forms an orthonormal basis in $L^{2}_{\phi}(\Omega)$. From [36, Section 3], one knows that the weighted heat kernel $H^{\phi}(x,y,t)$ on $\Omega$ has the following eigenfunction expansion formula
@@
In fact, the formula (4.1) was derived in [36, Section 3] by using several properties of the heat semigroup generated by the Witten-Laplacian $\Delta_{\phi}$. One can also check our recent work [67, Section 5] for a simpler way to get the formula (4.1) and also several other properties of the weighted heat kernel.
@@
Similar to what we have pointed out in [47, Remark 6.9], one cannot get the rigidity characterization for the equality in Cheng-type eigenvalue comparisons (1.14) or (1.15) through the approach used in this section. In fact, if $\lambda_{1,\phi}(B(q,r_{0}))=\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ here, we can only get $\lim_{t\rightarrow\infty}H^{\phi}(q,q,t)=\lim_{t\rightarrow\infty}H^{\phi}_{-}(0,t)$. We are not sure whether there exists some $t_{0}$ such that $H^{\phi}(q,q,t_{0})=H^{\phi}_{-}(0,t_{0})$ or not, which leads to the fact that we cannot use the rigidity characterization for the weighted heat kernel comparison (1.7) in Theorem 1.6. Same story happens to $\lambda_{1,\phi}(B(q,r_{0}))=\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right)$. By the way, the isometric rigidity characterizations for the eigenvalue comparisons (1.14) and (1.15) in Theorems 1.17-1.18 will be proven in Appendix B using another approach.

### v2

@@ 2 Preliminaries and some useful facts
and so $\sqrt{|g|}=\det\mathbb{A}(s,\xi)$. Therefore, by applying (2.2) and (2.3), it follows that the volume $\mathrm{vol}(B(q,r))$ of the geodesic ball $B(q,r)$ on $M^{n}$ is given by
@@
One has $\ell(q)=\max_{\xi\in S_{q}M^{n}}d_{\xi}$ (see page 704 of [33, Section 2] for the explanation). Clearly,
@@
As pointed out in [33, Remark 2.4] or [49, Remark 2.5], since the radial distance is $r(x)=d_{M^{n}}(x,q)$ for $x=\gamma_{\xi}(s)=\exp_{q}(s\xi)$, the parameter $s$ may be seen as the argument of the continuous function $\kappa:[0,l)\rightarrow\mathbb{R}$ in Definition 2.2 or Definition 2.3. Therefore, one has $\frac{d}{ds}|_{x}=\nabla r(x)=v_{x}$, which implies (2.12) and (2.13) become
@@
Similarly, one can define the concept that the radial Ricci/sectional curvature has an upper/lower bound (with respect to some point) which is given by a continuous function of the radial distance parameter. By the way, the concept that the radial (Ricci or sectional) curvature is bounded at some point has been improved to a more general setting, i.e. the integral radial (Ricci or sectional) curvature is bounded with respect to some point – see [53, Definitions 2.1 and 2.2]. In fact, the quantities defined in [53, Definitions 2.1 and 2.2] can be used to measure the deviation of the radial (Ricci or sectional) curvature (with respect to some point $q\in M^{n}$) from a prescribed continuous function $(n-1)\kappa(r(x))$ or $\kappa(r(x))$, where $r(x)=d_{M^{n}}(x,q)$ denotes the intrinsic distance to $q$ on the Riemannian $n$-manifold $M^{n}$.
Now, we wish to point out an important truth, that is, for a prescribed $n$-dimensional complete manifold $M^{n}$, one can always construct the optimal continuous functions $\kappa_{\pm}(q,s)$ with respect to a point $q\in M^{n}$, satisfying Definitions 2.2 and 2.3 respectively. This truth has already been shown in our previous works [33, p. 706], [49, pp. 378-379]. However, for convenience to readers, we wish to repeat the construction here. Recall that for $\xi\in S_{q}^{n-1}\subseteq T_{q}M^{n}$, $\gamma_{\xi}(s)=\exp_{q}(s\xi)$ and its derivative $\gamma^{\prime}_{\xi}(s)$ depend smoothly on the variables $(s,\xi)$. Set $\mathbb{D}_{q}:=\{(s,\xi)\in[0,\infty)\times S_{q}^{n-1}|0\leq s<d_{\xi}\}$ with closure $\overline{\mathbb{D}}_{q}:=\{(s,\xi)\in[0,\infty)\times S_{q}^{n-1}|0\leq s\leq d_{\xi}\}$. Then we can define
@@
As explained in [33, pp. 706-707] or [49, p. 379], by using the properties of warped product manifolds, one knows that the radial sectional curvature, and the radial component of the Ricci tensor of the spherically symmetric manifold $0,l)\times_{f(s)}\mathbb{S}^{n-1}$ with the base point $q$ are given by
@@
Clearly, if $r=0$ and $R=r_{0}$, then the annulus $A^{\Gamma}_{0,r_{0}}(q)$ would be exactly the geodesic ball $B(q,r_{0})$ on the given Riemannian $n$-manifold $M^{n}$. In [51, Theorem 3.2] (resp., [51, Theorem 4.3]), we have successfully obtained a relative volume comparison for the annulus $A^{\Gamma}_{r,R}(q)$ on a complete Riemannian $n$-manifold $M^{n}$ having a radial Ricci curvature lower bound $(n-1)f^{\prime\prime}(s)/f(s)$ with respect to the point $q\in M^{n}$ (resp., a radial sectional curvature upper bound $f^{\prime\prime}(s)/f(s)$ with respect to the point $q$). Moreover, those two volume comparison results would immediately imply our Bishop-type volume comparisons (i.e. Theorems 2.6 and 2.7 here) by suitably choosing radius parameters for annuluses – see [51, Corollaries 3.4 and 4.4] for details. In 2023, we attempted to improve our previous volume comparison results mentioned here to the setting that the given complete Riemannian $n$-manifold has integral radial curvature bounds (i.e. a more weaker curvature assumption), and this attempt was successful – see [53, Section 3] for details.
@@
In [31, Section 3], we have given several lower bounds for the first eigenvalue of weighted $p$-Laplacian on submanifolds with locally bounded weighted mean curvature.
@@ 3 Weighted heat kernel comparison theorems
This question is crucial for the eigenvalue problem (1.11) and the answer is affirmative. In fact, as mentioned in A. Grigor’yan’s survey [38, Section 2], by suitably making adjustments to those related arguments in e.g. [19, 27, 34, 65], it is not hard to get the following result.
@@
([38, Section 2]) The operator $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a self-adjoint nonpositive definite operator in $L^{2}_{\phi}$, where the space $W^{2,2}_{0,\phi}(\Omega)$ is defined as $W^{2,2}_{0,\phi}(\Omega):=\{\rho\in W^{1,2}_{0,\phi}(\Omega)|\Delta_{\phi}\rho\in L^{2}_{\phi}(\Omega)\}$. Moreover, the operator $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a unique self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$ with the domain in $W^{1,2}_{0,\phi}(\Omega)$. If in addition $\Omega$ is geodesically complete, then the operator $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$ is essentially self-adjoint, i.e., $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a unique self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$.
Once we have Theorem 3.1 for the extension of the operator $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$. The spectrum of the eigenvalue problem (1.11) is actually the spectral structure of the self-adjoint operator $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ in $L^{2}_{\phi}(\Omega)$, which is only discrete and can be deduced from the compactness of the resolvent $\left(-\Delta_{\phi}+I\right)^{-1}$, together with the spectral theory of compact self-adjoint differential operators. See e.g. [38, Subsection 2.2] for details. Combining with the variational principle, one can get two basic facts for the eigenvalue problem (1.11) listed in Section 1.
As explained in [38, Section 3], once the self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$ has been set up as in Theorem 3.1, the heat semigroup $e^{t\Delta_{\phi}}$ related to the Witten-Laplacian $\Delta_{\phi}$ can be well-defined. Then one can get the following properties for the weighted heat kernel by using the tool of heat semigroup.
@@
(see e.g. [38, Theorem 3.3 in Section 3]) Let $M$ be a complete Riemannian manifold and $\phi$ be a real-valued smooth function on $M$. Denote by $\mathbb{R}^{+}$ the set of all positive real numbers. Then there exists a weighted heat kernel $H^{\phi}(x,y,t)\in C^{\infty}(M\times M\times\mathbb{R}^{+})$ such that
@@
It is easy to see that (1)-(2) in Lemma 3.2 have close relation with the existence of the weight heat kernel. In [38, Theorem 3.3 in Section 3], the property (3) is the symmetry of the weight heat kernel, while the property (4) is the semigroup identity of the weight heat kernel. In [38, Section 3], A. Grigor’yan mentioned at least two approaches to prove the existence of the weight heat kernel — the classical one is the usage of a parametrix of the heat equation, and the other one is the usage of eigenfunction expansion formula in an exhaustion of relatively compact open sets $\{\Omega_{i}\}_{i=1}^{\infty}$ of $M$. For details, one can check e.g. [11, 16, 19, 27, 65] as references.
@@
It is easy to check that $\mathcal{L}^{\phi}$ is uniformly parabolic, and then the assertion in Theorem 3.1 directly follows from the standard theory of second-order parabolic partial differential equations (see e.g. [16, p. 180] or [32, Chapter 7]). ∎
@@
If the potential function $\phi$ has Property 1, then the weighted heat kernel $H_{+}^{\phi}$ (resp, $H^{\phi}_{-}$) on a geodesic ball $\mathscr{B}_{n}(q^{+},\cdot)$ (resp, $\mathscr{B}_{n}(q^{-},\cdot)$) of the model space $M^{+}$ (resp, $M^{-}$) depends on the space variable $r_{1}$ (resp., $r_{2}$) can be observed by the decomposition of the Witten-Laplacian on $\mathscr{B}_{n}(q^{+},\cdot)$ (resp, $\mathscr{B}_{n}(q^{-},\cdot)$) in geodesic spherical coordinates at $q^{+}$ (resp, $q^{-}$) — see the $3^{\rm{rd}}$ line in the proof of Lemma 6.1 for the detailed expression. The monotonicity of $H_{+}^{\phi}$ (resp, $H^{\phi}_{-}$) with respect to the time variable can be obtained by using an almost the same argument to that in the proof of [19, Lemma 2.3]. But there are two differences we wish to point out as follows:
@@
    where $a_{i}(r_{1})$ are smooth functions on $M^{+}$. This directly implies the expression (2.11) in [19]. In the weighted case, [38, Theorem 3.9] tells us that for sufficiently small $0<r_{1},t<\epsilon$, there exists a smooth positive function $u(r_{1})$ such that the weighted heat kernel $H_{+}^{\phi}(r_{1},t)$ satisfies
@@
    which leads to the truth that the analysis process in the proof of [19, Lemma 2.3] before the equation (2.20) can be transplanted to the non-weighted case without any big obstacle. Same story happens to $H^{\phi}_{-}(r_{2},t)$. By the way, a detailed proof of the above asymptotic expansion for the weighted heat kernel can be found in [10, 36].
@@
We wish to use a similar argument to that in the proof of our previous work [49, Theorem 6.6].
@@
$\qquad H^{\phi}(q,y,t)-H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)$ (3.5)
@@
$\geq 0,$
where the fact $d\theta(s,\xi)/ds=(n-1)\left[J(s,\xi)/f(s)\right]^{n-2}\left[J(s,\xi)/f(s)\right]^{\prime}\leq 0$ shown in Theorem 2.6 has also been used. This directly implies $H^{\phi}(q,y,t)\geq H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)$, which is exactly the weighted heat kernel comparison (1.7). When the equality in (1.7) holds at some $(y_{0},t_{0})\in B(q,r_{0})\times(0,\infty)$, by Theorem 3.4 one has $H^{\phi}(q,y,t)=H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)=H^{\phi}(q,y_{0},t_{0})$ on $B(q,r_{0})\times[0,t_{0}]$. Together with (3.5), we know that
@@ 4 Applications in spectral geometry
Theorem 3.1 tells us that if $\Omega\subseteq M^{n}$ is geodesically complete, with $M^{n}$ a given complete Riemannian $n$-manifold, then $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a unique self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$, and this fact can be used to prove that the eigenvalue problem (1.11) only has discrete spectrum, which consists of eigenvalues given in the sequence (1.12). Let $\psi_{i}$ be an eigenfunction corresponding to the $i$-th Dirichlet eigenvalue $\lambda_{i,\phi}(\Omega)$ such that $\{\psi_{i}\}_{i=1}^{\infty}$ forms an orthonormal basis in $L^{2}_{\phi}(\Omega)$. From [38, Section 3], one knows that the weighted heat kernel $H^{\phi}(x,y,t)$ on $\Omega$ has the following eigenfunction expansion formula
@@
In fact, the formula (4.1) was derived in [38, Section 3] by using several properties of the heat semigroup generated by the Witten-Laplacian $\Delta_{\phi}$. One can also check our recent work [72, Section 5] for a simpler way to get the formula (4.1) and also several other properties of the weighted heat kernel.
@@
Similar to what we have pointed out in [49, Remark 6.9], one cannot get the rigidity characterization for the equality in Cheng-type eigenvalue comparisons (1.14) or (1.15) through the approach used in this section. In fact, if $\lambda_{1,\phi}(B(q,r_{0}))=\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ here, we can only get $\lim_{t\rightarrow\infty}H^{\phi}(q,q,t)=\lim_{t\rightarrow\infty}H^{\phi}_{-}(0,t)$. We are not sure whether there exists some $t_{0}$ such that $H^{\phi}(q,q,t_{0})=H^{\phi}_{-}(0,t_{0})$ or not, which leads to the fact that we cannot use the rigidity characterization for the weighted heat kernel comparison (1.7) in Theorem 1.6. Same story happens to $\lambda_{1,\phi}(B(q,r_{0}))=\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right)$. By the way, the isometric rigidity characterizations for the eigenvalue comparisons (1.14) and (1.15) in Theorems 1.17-1.18 will be proven in Appendix B using another approach.